Compute the sum Σ from k=1 to infinity of (-7/6)^(k-1)
The correct answer is one of the following. Which one is correct?
A) 7/6
B) 7
C) 7/13
D) does not exist

Answer by mccravyedwin(407)


That's an infinite geometric series with common ratio r = -7/6.
But an infinite geometric series diverges (does not exist) when |r| > 1.

But |-7/6| = 7/6 > 1, so the series does not exist.

Edwin